# High-performance Vector-length Agnostic Quantum Circuit Simulations on ARM Processors

### v2

@@ IV Vector-length Agnostic Quantum Simulations
Figure 5 illustrates an example of applying the Hadmard gate on the 3rd qubit for an $n$-qubit system.When assigning the state groups to threads, we iterate the 3rd bit 0/1 and keep the other bits the same within a group. The number of states in a parallel group depends on the gate size, e.g., a single-qubit gate would have two states. Extending the case of a single-qubit gate to a k-qubit gate, the number of states per group is $2^{k}$. Considering the vectorization, in each thread, we can calculate $numVals$ state groups simultaneously using VLA vectorization. Here, $numVals$ is calculated as $\frac{Vector\;Length(VLEN)}{Element\;Width(ELEN)}$, where $VLEN$ is an architecture characteristic and $ELEN$ is usually 32-bit for single-precision and 64-bit for double-precision. $numVals$ represents the maximum number of data elements that can be operated together in one vector register.
@@ IV-A VLEN-adaptive Memory Access Optimization
In this work, we choose to adjust the memory layout in place during initialization and maintain it throughout the gate applying stage. We adjust to continuously store $numVals$ real floats followed by $numVals$ imaginary floats in memory to ensure that consecutive addresses are loaded from memory into the vector register. Compared to a VLS implementation, for a VLA design, we need to ensure $numVals$ is adaptive to different vector lengths on different architectures. Step 1 of Figure 5 illustrates the memory layout before and after adjustment of the first 16 states when $numVals$ is equal to 4. Assuming a gate is applied to the $i$-th qubit $q_{i}$, where $i$ is greater than $log_{2}(numVals)$, it will be consecutively loaded into the vector registers for real and imaginary parts, and then $numVals$ state groups are computed simultaneously. This VLEN-adaptive memory layout provides more efficient memory access for VLA architectures with requiring only two additional loops out of the size $2^{n-1}$ in the initialization stage for memory adjustment.
@@ IV-C Fine-grained Loop Control
One challenge in efficient vectorization is to identify loops that will improve performance by vectorization, but avoid those that will degrade performance. Ideally, with vectorization, we can reduce the size of loops by $numVals$ times. When applying gates on qubit $q_{i}$, where $2^{i}$ is larger than $numVals$, no predicate elements are needed, i.e., they are regular loops shown in Figure 5. However, if $2^{i}$ is smaller than $numVals$, they become irregular loops. The two consecutive states are included in the same state group in irregular loops, but only the state in different groups can be processed in parallel. Using the predicate in Figure 1, we process these states in sequence. We still load $numVals$ consecutive values, but mask out unnecessary states one by one. When vectorizing an irregular loop, if there are more than two active elements in the predicate register, vectorization may still gain performance. However, if the predicate enables only a single active element, it is more efficient to use the scalar unit. We analyze all gates and find the only special condition is multi-qubit gates acting on lower qubits ($q_{1}$ to $q_{log_{2}(numVals)}$), will process only one active element per vector instruction.
@@ V Implementation
For loops with the non-uniform strides in irregular memory access, as shown in Figure 3, explicit porting with predicate settings is needed. Listing 2 illustrates the algorithm to setting predicate register with the time complexity $O(L)$, where L represents the count of gate-applied qubits lower than $numVals$. qs stores indices of acted qubits. The idex generates a sequence of sequences with a step size of 1 starting from 0 and then it operates bitwise AND with the bitmask tmp to generate the value of predicate pg.
@@ VI Experimental Setup
TABLE III: The number of gate operations on qubit $q_{i}$ in a $N$-qubit system. numValsnumVals is the number of elements in a vector instruction.
@@ VII-A Vectorization Activities
While the SVE version achieves high AVL, close to their theoretical limit of $numVals$, the end-to-end performance speedup is much lower. For instance, QFT gets 12.6 AVL but only $4.4\times$ speedup on A64FX. Therefore, we apply the top-down method [40] to identify the performance bottlenecks. Table IV reports the percentage of frontend stall cycles and backend stall cycles over the total stall cycles as FE and BE, respectively. Across five benchmarks, frontend stall is merely 0.09% - 0.3% and 4.9% - 14.6%, while the backend stall reaches 99.7%-99.9% and 85.4%-95.1% on Grace and A64FX. Therefore, we further analyzing backend stalls, and find that backend memory stalls occupy about 39.4%-48.6%. Since backend_boundbackend\_bound includes core_boundcore\_bound waiting for drivers and execution ports like ALU/FPU, and also memory_boundmemory\_bound waiting for memory load/store, the results indicate SVE design to be balanced between memory-bound and core-bound.
@@ XI-A Computational Artifact $A_{1}$
Our implementation can be performed on ARM CPU equipped with SVE. The memory capability requires more than $2^{(N+3)}$ B for $N$-qubit simulation. The experiments were performed on Nvidia Grace, Graviton and A64FX CPU.

### v3

@@ IV Vector-length Agnostic Quantum Simulations
Figure 5 illustrates an example of applying the Hadmard gate on the 3rd qubit for an $n$-qubit system.When assigning the state groups to threads, we iterate the 3rd bit 0/1 and keep the other bits the same within a group. The number of states in a parallel group depends on the gate size, e.g., a single-qubit gate would have two states. Extending the case of a single-qubit gate to a k-qubit gate, the number of states per group is $2^{k}$. Considering the vectorization, in each thread, we can calculate numValsnumVals state groups simultaneously using VLA vectorization. Here, numValsnumVals is calculated as $\frac{Vector\;Length(VLEN)}{Element\;Width(ELEN)}$, where VLENVLEN is an architecture characteristic and ELENELEN is usually 32-bit for single-precision and 64-bit for double-precision. numValsnumVals represents the maximum number of data elements that can be operated together in one vector register.
@@ IV-A VLEN-adaptive Memory Access Optimization
In this work, we choose to adjust the memory layout in place during initialization and maintain it throughout the gate applying stage. We adjust to continuously store numValsnumVals real floats followed by numValsnumVals imaginary floats in memory to ensure that consecutive addresses are loaded from memory into the vector register. Compared to a VLS implementation, for a VLA design, we need to ensure numValsnumVals is adaptive to different vector lengths on different architectures. Step 1 of Figure 5 illustrates the memory layout before and after adjustment of the first 16 states when numValsnumVals is equal to 4. Assuming a gate is applied to the $i$-th qubit $q_{i}$, where $i$ is greater than $log_{2}(numVals)$, it will be consecutively loaded into the vector registers for real and imaginary parts, and then numValsnumVals state groups are computed simultaneously. This VLEN-adaptive memory layout provides more efficient memory access for VLA architectures with requiring only two additional loops out of the size $2^{n-1}$ in the initialization stage for memory adjustment.
@@ IV-C Fine-grained Loop Control
One challenge in efficient vectorization is to identify loops that will improve performance by vectorization, but avoid those that will degrade performance. Ideally, with vectorization, we can reduce the size of loops by numValsnumVals times. When applying gates on qubit $q_{i}$, where $2^{i}$ is larger than numValsnumVals, no predicate elements are needed, i.e., they are regular loops shown in Figure 5. However, if $2^{i}$ is smaller than numValsnumVals, they become irregular loops. The two consecutive states are included in the same state group in irregular loops, but only the state in different groups can be processed in parallel. Using the predicate in Figure 1, we process these states in sequence. We still load numValsnumVals consecutive values, but mask out unnecessary states one by one. When vectorizing an irregular loop, if there are more than two active elements in the predicate register, vectorization may still gain performance. However, if the predicate enables only a single active element, it is more efficient to use the scalar unit. We analyze all gates and find the only special condition is multi-qubit gates acting on lower qubits ($q_{1}$ to $q_{log_{2}(numVals)}$), will process only one active element per vector instruction.
@@ V Implementation
For loops with the non-uniform strides in irregular memory access, as shown in Figure 3, explicit porting with predicate settings is needed. Listing 2 illustrates the algorithm to setting predicate register with the time complexity $O(L)$, where L represents the count of gate-applied qubits lower than numValsnumVals. qs stores indices of acted qubits. The idex generates a sequence of sequences with a step size of 1 starting from 0 and then it operates bitwise AND with the bitmask tmp to generate the value of predicate pg.
@@ VI Experimental Setup
TABLE III: The number of gate operations on qubit $q_{i}$ in a NN-qubit system. numValsnumVals is the number of elements in a vector instruction.
@@ VII-A Vectorization Activities
While the SVE version achieves high AVL, close to their theoretical limit of numValsnumVals, the end-to-end performance speedup is much lower. For instance, QFT gets 12.6 AVL but only $4.4\times$ speedup on A64FX. Therefore, we apply the top-down method [40] to identify the performance bottlenecks. Table IV reports the percentage of frontend stall cycles and backend stall cycles over the total stall cycles as FE and BE, respectively. Across five benchmarks, frontend stall is merely 0.09% - 0.3% and 4.9% - 14.6%, while the backend stall reaches 99.7%-99.9% and 85.4%-95.1% on Grace and A64FX. Therefore, we further analyzing backend stalls, and find that backend memory stalls occupy about 39.4%-48.6%. Since backend_boundbackend\_bound includes core_boundcore\_bound waiting for drivers and execution ports like ALU/FPU, and also memory_boundmemory\_bound waiting for memory load/store, the results indicate SVE design to be balanced between memory-bound and core-bound.
@@ XI-A Computational Artifact $A_{1}$
Our implementation can be performed on ARM CPU equipped with SVE. The memory capability requires more than $2^{(N+3)}$ B for NN-qubit simulation. The experiments were performed on Nvidia Grace, Graviton and A64FX CPU.